OpenStudy (brinazarski):

Find the product of the nth root of 1. I got either one or infinity... of course I'm not entirely sure what the question is asking...

OpenStudy (brinazarski):

... you lost me at k ._.

OpenStudy (brinazarski):

So is it one...? I don't understand a word on that page, and don't think I learned about roots of unity...

OpenStudy (brinazarski):

I don't know what math this is... I was just placed in my school's Math Team and have to do the homework, but so far I've missed about 2-3 weeks of class since I was just put in it...

OpenStudy (brinazarski):

"In mathematics, a root of unity, or de Moivre number, is any complex number that equals 1 when raised to some integer power n."-Wikipedia. So all I understand from this is that the answer is one...?

OpenStudy (brinazarski):

I sucked at Algebra I, aced Geometry, was put into an advanced Trig and Algebra II class and got a 76, but then I became good at Algebra I... on Regents level I got an 89 (in Trig/AII, in A I I'm now almost at a 100 on Regents level I think). So that's the level I'm at when it comes to math.

OpenStudy (brinazarski):

Yes... and specifically the New York system lol

OpenStudy (brinazarski):

Basically, I'm great at Algebra I and Geometry, okay at Trig and Algebra II, and I haven't done any math beyond that. That make any more sense?

OpenStudy (brinazarski):

I'm 17

OpenStudy (brinazarski):

Last year of high school

OpenStudy (brinazarski):

Ah... so, I'm looking at the equation. I don't know n or k, but you told me that k is the roots of unity, which is anything that equals 1 when raised to n... would that make n = 0? O_o

OpenStudy (brinazarski):

I don't understand the "by induction" part that the person is asking on the non-Wiki page..

OpenStudy (brinazarski):

Wait

OpenStudy (brinazarski):

Are the roots of unity imaginary?

hartnn (hartnn):

i will try to keep this as simple as possible: let x be the variable to represent n'th root of 1 \(x=1^{(1/n)}\) raising to n'th power \(x^n=1 \\ x^n-1=0\) is your polynomial equation. Now we use a property that the product of roots of general equation \(Ax^n+Bx^{n-1}+Cx^{n-2}+....N=0\) is \((-1)^n\) times the constant co-efficient. So the product of roots of x^n-1 = 0 will be \((-1)^n (-1)=(-1)^{n+1}\) this is your product of n'th root of 1. hope this makes some sense.Ask if there is doubt in any step.

OpenStudy (brinazarski):

Is that the answer then, or is there more that needs to be solved?

hartnn (hartnn):

that is the answer (-1)^{n+1}

OpenStudy (brinazarski):

I'm a little lost at the property of the product of the roots... where did you get (-1)^n? Otherwise, I think I understand the rest.

hartnn (hartnn):

its a general formula , it can be proved, but i donno the best way to prove it. the property is listed here, where u have sum of roots, product of roots and other formulas....just go through it.... http://en.wikipedia.org/wiki/Vieta%27s_formulas

OpenStudy (brinazarski):

Okay! Thank you very much! You have been so helpful today! :)

hartnn (hartnn):

welcome very much \(\ddot\smile\)